Two cars, A and B, having average speeds of 30 and 40 mph, respectively are 280 miles apart they start moving towards each other at 3pm, At what time will they meet?
Answer by checkley77(12844)

30x+40x=280
70x=280
x=280/70
x=4 hours.
3+4=7 pm. they will meet.